OpenStudy (anonymous):

the perimeter of a retangular room is 166ft.find the length and width of the room if the length is 5 ft longer than twice the width.

OpenStudy (anonymous):

26x57

OpenStudy (anonymous):

P = 2w + 2l 166 = 2w + 2(2w+5) 166 = 6w + 10 156 = 6w 26 = w 52 + 5 = 57 = l

OpenStudy (anonymous):

let the width be =w length=2(w+5) 166=2w+2(2w+5) 166=2w+4w+10 156=6w w=26 l=2(26)+5 57